OpenStudy (kenkaneki12280000):

The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –3 6 3 4 Line 2 x y –6 7 9 2 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (kenkaneki12280000):

@soprano.h.d0816

OpenStudy (kenkaneki12280000):

@soprano.h.d0816

OpenStudy (texaschic101):

line 1....slope = -1/3....y int = 5 line 2...slope = - 1/3...y int = 5 infinite solutions

OpenStudy (kenkaneki12280000):

@texaschic101 i dont know how to give medals but i give u best response

OpenStudy (texaschic101):

thats exactly how to give medals.....thx

OpenStudy (kenkaneki12280000):

oh an u help meh with another one

OpenStudy (texaschic101):

sorry...i am already running late....i will be back on later

OpenStudy (kenkaneki12280000):

ok thx for your help

OpenStudy (texaschic101):

yw

OpenStudy (kenkaneki12280000):

does that mean yeah whatever

OpenStudy (texaschic101):

lol...no, it simply means your welcome